In December 2016, Custom Mfg. established its predetermined overhead rate for jobs produced during 2017 by using the following cost predictions: overhead costs, $750,000, and direct materials costs $625,000. At year-end 2017, the companys records show that actual overhead costs for the year are $830,000. Actual direct material cost had been assigned to jobs as follows. Jobs completed and sold Jobs in finished goods inventory Jobs in work in process inventory Total actual direct materials cost $513, 750 102, 750 68,500 $685, ΘΘΘ 1. Determine the predetermined overhead rate for 2017. 2&3. Enter the overhead costs incurred and the amounts applied during the year using the predetermined overhead rate and determine whether overhead is overapplied or underapplied 4. Prepare the adjusting entry to allocate any over- or underapplied overhead to Cost of Goods Sold. Complete this question by entering your answers in the tabs below Prev 1 of 7Next

media%2F069%2F069f8278-5174-4dde-99e5-d8
Answer #1
1
Overhead Rate
Choose Numerator: Choose Denominator: Overhead Rate
Estimated overhead costs / Estimated direct materials costs = Overhead Rate
750000 / 625000 = 120%
2&3
Factory Overhead
Actual overhead 830000 822000 Applied overhead
Underapplied overhead 8000
4
Dec.31 Cost of goods sold 8000
        Factory overhead 8000
Note: Applied overhead =685000*120% = $822000